OpenStudy (anonymous):

The sum of two consecutive odd integers is 244. What is the smaller integer? 119 125 123 121

OpenStudy (mathmate):

I'll set it up for you. Let x=the smaller odd integer. x+(x+2)=244 solve for x.

OpenStudy (anonymous):

it's 121

OpenStudy (mathmate):

Correct, well done!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Let x = the first odd integer then, x + 2 is the next (consecutive) odd integer x + (x + 2) = 244 [the sum of two consecutive odd integers] solve for x 2x + 2 = 244 2x = 242 x = 121 121 + 123 = 244

OpenStudy (mathmate):

You're welcome!

OpenStudy (anonymous):

first you should solve for x ok as soon you do that you will find 121 which is "d"